The normal force of an object on the inclined plane is 6.5 N and the mass of the object is 0.80 kg. Find


a) the weight of the object b) the frictional force to slide the object c) What is the coefficient of


friction on the inclined plane ?



1
Expert's answer
2021-04-13T06:42:07-0400

1) The weight of the object is "P = m g \approx 7.85 N"

2) Writing the projections of the forces, onto axes,along and perpendicular to the incline, obtain "m g \sin \alpha = F_f", "N = m g \cos \alpha".

From the second equation, "\cos \alpha = \frac{N}{m g}", hence from the first equation:"F_f = m g \sin \alpha = m g \sqrt{1 - \cos^2 \alpha} = m g \sqrt{1 - \left( \frac{N}{m g} \right)^2} \approx 4.4 N"

3) Since "F_f = \mu N", coefficient of friction is "\mu = \frac{F_f}{N} \approx 0.68".